hi please help me solve this problem ( solve the linear system using substitution) 3y+x=2, y-x=-6
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

3y+x=2
y-x=-6
---
Solve the 2nd for "y":
y = x-6
---
Substitute into the 1st and solve for "x":
3(x-6)+x = 2
3x-18 + x = 2
4x = 20
x = 5
----
Solve for "y":
y = x-6
y = 5-6
y = -1
=============
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi
solving the linear system using substitution
y-x= -6 | x = (y+6)
3y+x=2
3y + (y+6) = 2 |substituting for x
solving for y
4y = -4
y = -1 x = 5 (y+6)
CHECKING our Answer***
y-x= -6
-1-5 = -6